John is twice as old as Peter. The sum of their ages is 36 years. In 5 years time, how old will John be?
Answer by TimothyLamb(4379)

j = 2p
j + p = 36
---
j + p = 36
2p + p = 36
3p = 36
p = 12
j = 24
---
Answer:
j + 5 = 29
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php